# String graphs are quasi-isometric to planar graphs

### v1

@@
Abstract
@@ 1 Introduction
Every planar graph is a string graph, although the structure of string graphs is much more complicated. For instance, while planar graphs can be recognized in linear time [47], recognition of string graphs is NP-hard [52, 73]. There are at most $c^{n}n!$ planar graphs on $n$ vertices [22, 65], while the number of string graphs on $n$ vertices is $2^{(\frac{3}{4}+o(1)){n\choose 2}}$ [68]. Every planar graph is 4-colourable [5, 71], however there are triangle-free string graphs with arbitrarily large chromatic number [69].
@@
In other words, string graphs are quasi-isometric (or even bi-Lipschitz equivalent) to planar graphs. For finite string graphs the proof of Theorem˜1.1 yields a polynomial time (with respect to the size of a given representation of $S$ in the input) algorithm for generating such a planar graph $G$. Unfortunately, are string graphs that require exponentially sized representations [53]. However, natural subclasses such as segment intersection graphs or more generally intersection graphs of convex compact path-connected subsets of $\mathbb{R}^{2}$, do have polynomially sized representations.
Theorem˜1.1 has numerous applications. We first discuss some direct corollaries for string graphs. Afterwards, we shall discuss two further quasi-isometry results on metric planar graphs and complete Riemannian planes.
@@
One of the most natural classes of graphs to consider is of course planar graphs. Ostrovskii and Rosenthal [66] proved that planar graphs have bounded asymptotic dimension. Fujiwara and Papasoglu [35] improved this bound to 3 and Jørgensen and Lang [49] proved that planar graphs have asymptotic dimension and furthermore Assouad–Nagata dimension at most 2, which is tight due to the infinite planar grid. Asymptotic dimension and Assouad–Nagata dimension are persevered by quasi-isometries. Thus, by Theorem˜1.1, we obtain the following.
@@
Bonamy, Bousquet, Esperet, Groenland, Liu, Pirot, and Scott [11] gave a proof that planar graphs have Assouad–Nagata dimension at most 2, which yields a polynomial time algorithm for computing the corresponding covers. Thus, the corresponding covers in Theorem˜1.2 can also be computed in polynomial time (in terms of the size of the provided representation of the string graph given in the input). For some algorithmic applications of such covers, see [7, 12, 13, 30, 48, 54, 56].
Theorem 1.3.
For a more extensive survey and discussion of applications of such embeddings, see [44].
@@
We now turn our attention to some further quasi-isometry results on metric planar graphs and complete Riemannian planes.
A metric graph $H$ is a graph along with an assignment of non-negative edge lengths (in this paper, the length of edges in (non-metric) graphs is otherwise always uniformly equal to 1). We can view metric graphs as a length space in which the edges are replaces by intervals of the corresponding lengths. When the edge lengths are within $D\subseteq(0,\infty)$, we say that $H$ is an $D$-metric graph.
@@
Beyond the general interest in metric planar graphs and also in groups acting on planar surfaces [39], one other motivation of Georgakopoulos and Papasoglu [42] was their coarse Kuratowski conjecture that graphs forbidding $K_{5}$ and $K_{3,3}$ as $K$-fat minors are quasi-isometric to planar graphs. Although Theorem˜1.6 and Theorem˜1.1 provide substantial progress towards this conjecture, we remark that the closely related graph minor analogue that graphs forbidding $H$ as a $K$-fat minor are quasi-isometric to $H$-minor-free graphs is false [2, 21], (although it is known to be true for some small graphs $H$ [3, 4, 16, 36, 42, 60, 64]). Such $K$-fat $H$-minor-free graphs need not even be quasi-isometric to $H^{\prime}$-minor-free graphs for any graph $H^{\prime}$ [2]. However, with Distel and Hickingbotham [18], we will use the main technical result (see Theorem˜5.1) of this paper to prove an approximate graph minor version of Theorem˜1.6 that metric $H$-minor-free graphs are quasi-isometric to $H^{\prime}$-minor-free graphs. We remark that the constants of the quasi-isometries necessarily depend on $H$ [20].
It is well known (see [11, 59, 72]) that complete Riemannian planes are quasi-isometric to locally finite [^2] A metric graph is locally finite if every bounded radius ball contains at most a finite number of vertices. metric planar graphs. This was used for instance to show that complete Riemannian planes have Assouad–Nagata dimension at most 2 [11] and in Maillot’s [59] aforementioned classification finitely generated groups virtually free groups as being those quasi-isometric to certain planar Riemannian surfaces. By Theorem˜1.6, we obtain the following refinement, which significantly generalizes the previously discussed fact that finitely generated groups quasi-isometric to complete Riemannian planes are quasi-isometric to planar graphs.
@@
In [17] we will use the main technical result (see Theorem˜5.1) of the present paper to extend Theorem˜1.7 to higher genus Riemannian surfaces and also refine the theorem so that the quasi-isometry from the graph to the Riemannian surface is a continuous embedding. Such a result does not hold for complete Riemannian surfaces of unbounded genus [17], (although, it does if we use metric graphs [11, 59, 72] instead of graphs).
@@
In Section˜2 we introduce basic definitions, deduce Theorem˜1.6 from Theorem˜1.1, and also reduce proving Theorem˜1.1 to its finite case (Theorem˜2.5). We also introduce the necessary framework for our approach to proving Theorem˜1.1, namely through viewing string graphs as region intersection graphs, and the notion of “impressions” which provides a convenient way of finding quasi-isometries from region intersection graphs. The rest of the proof in Sections˜3, 4 and 5 is essentially a delicate technical induction. In Section˜3 we prove a weakening of Theorem˜1.1 for the subclass outerstring graphs (Theorem˜3.1). This is used in Section˜4 to find a technical structure tailored for use in the inductive argument that we call an “encasing” (see Lemma˜4.4). Finally, in Section˜5, we prove Theorem˜1.1 using Lemma˜4.4.
@@ 2 Preliminaries
Proof.
Let $f:\mathcal{H}\to\mathcal{M}$ be chosen so that $f(H)\in I_{\mathcal{M}}(H)$ for all $H\in\mathcal{H}$.
@@
Next we show that Theorem˜1.1 implies Theorem˜1.6. This is immediate from the following lemma.
@@
Proof.
Consider $H$ embedded in the plane. For each $u\in V(H)$, let $S_{u}$ be the subset of the plane consisting of all points of $H$ at distance at most 1 in $H$ from $u$. Let $S$ be the string graph with vertex set $V(H)$ in which each $u\in V(S)=V(H)$ is represented by $S_{u}$. Consider $u,v\in V(H)$ and let $P$ be a geodesic path from $u$ to $v$ in $H$. Choose vertices $x_{0},\ldots,x_{t}$ of $H$ along $P$ so that $x_{0}=u$, $x_{t}=v$ and for each $0\leqslant i<t$, the distance along $P$ between $x_{i}$ and $x_{i+1}$ is in $(1,2]$ (or if $d_{H}(u,v)\leqslant 1$, we just take $x_{0}=u$, $x_{1}=v$). So, $\lceil\frac{1}{2}d_{H}(u,v)\rceil\leqslant t\leqslant\lfloor d_{H}(u,v)\rfloor$. The lemma now follows since $x_{i}$ is adjacent to $x_{i+1}$ for each $0\leqslant i<t$. ∎
@@ 3 Outerstring graphs
Outerstring graphs are equivalent to region intersection graphs $\textbf{RIG}(G,\mathcal{H})$ for which $G$ is planar and every $H\in\mathcal{H}$ contains a vertex of the outer face of $G$. They can be thought of as an “outerplanar” analogue of string graphs, and are thus much simpler. The main aim of this section is to prove the following theorem on finding certain restricted impressions of such region intersection graphs. This will be crucial to our arguments on string graphs (or strictly speaking, Lemma˜3.4 and Lemma˜3.2 which easilly imply the following theorem will be).
@@
Figure 1: The graph $F_{6}$.
By Proposition˜2.1, Theorem˜3.1 implies that outerstring graphs are quasi-isometric to outerplanar graphs. This result is already easily implied by a theorem that graphs forbidding $K_{2,3}$ as a fat minor are quasi-isometric to cacti [16, 36], where a cacti is a planar graph in which every edge is on the outer face, or equivalently, a graph in which each maximal 2-connected subgraph is either an edge of a cycle. One might hope that Theorem˜3.1 could be improved so that $\textbf{IM}(G,\mathcal{M})$ is also a cacti. However this is not possible. For every $\ell\geqslant 3$, let $F_{\ell}$ be the outerplanar graph consisting of a cycle of length $\ell$ where for every edge of this cycle, we identify the edge with an edge of another cycle of length $\ell$ (see Figure˜1 for an illustration of the graph $F_{6}$). Then it can easily be seen that there is no $x,y$ such that every $(F_{\ell},\{\{u,v\}:uv\in E(F_{\ell})\})$ has a $(x,y)$-impression $(F_{\ell},\mathcal{M}_{\ell})$ in which $\textbf{IM}(F_{\ell},\mathcal{M}_{\ell})$ is a cacti. It is however a nice exercise to show that outerplanar graphs are quasi-isometric to cacti. With this, Theorem˜3.1 and Proposition˜2.1 then yeild another proof that outerstring graphs are quasi-isometric to cacti.
@@
Proof.
Clearly it is enough to handle the case that $G$ is connected. Choose some $H_{0}\in\mathcal{H}$ and for every non-negative integer $i$, let $L_{i}$ be the vertices of $\textbf{RIG}(G,\mathcal{H})$ at distance $i$ from $H_{0}$ in $\textbf{RIG}(G,\mathcal{H})$. For each $i$ we define the following two vertex sets in $G$. Let $D_{i}=\bigcup_{H\in L_{i}}H$, and let $A_{i}=D_{i}\backslash D_{i-1}$ (where $D_{-1}=\emptyset$). For each $A_{i}$, let $\mathcal{M}_{i}$ be the maximal connected vertex sets of $G[A_{i}]$. Let $\mathcal{M}=\bigcup_{i=0}^{\infty}\mathcal{M}_{i}$. See Figure˜2 for an illustration. We must show that $(G,\mathcal{M})$ is a $(11,9)$-impression of $(G,\mathcal{H})$.
Figure 2: An illustration of the resulting collection of disjoint connected sets $\mathcal{M}$. The connected sets in $\mathcal{H}$ are the red sets and the sets in $\mathcal{M}$ are the blue ones. The dashed purple lines indicate the partition of $\mathcal{M}$ into $\bigcup_{i=0}^{\infty}\mathcal{M}_{i}$.
@@
Lastly, we can now assume that $K$ satisfies Item˜4. Let $B_{1},B_{2},F$ be as in Item˜4. Note that the intersection of $K$ with the vertices of the outer face of $G$ induces a connected graph since $B_{1}$ and $B_{2}$ both contain a vertex on the outer face of $G$ and the face $F$ (which is distinct from the outer face). Let $Q$ be the vertex set of this connected induced subgraph. Let $p_{1}p_{2}\ldots p_{t}$ be the walk in $G[Q]$ along the outer face of $G$ (and containing all of) $Q$, where $p_{1}$ touches $B_{1}$ and $p_{t}$ touches $B_{2}$. Choose $1\leqslant\ell\leqslant t$ maximum such that that there exists a path with vertex set $L$ between $p_{\ell}$ and $N_{G}(B_{1})$ such that $L$ is contained in $K\cap(H_{\ell}\cup H^{\prime}_{\ell})$ for some intersecting (but not necessarily distinct) $H_{\ell},H^{\prime}_{\ell}\in\mathcal{H}$. Similarly, choose $1\leqslant r\leqslant t$ minimum such that that there exists a path with vertex set $R$ between $p_{r}$ and $N_{G}(B_{2})$ such that $R$ is contained in $K\cap(H_{r}\cup H^{\prime}_{r})$ for some intersecting (but not necessarily distinct) $H_{r},H^{\prime}_{r}\in\mathcal{H}$. Note that such $\ell,r$ exist since $K$ touches exactly $B_{1},B_{2}$ from $\mathcal{B}$, and $K$ (and every $H\in\mathcal{H}$) contains a vertex incident to the outer face of $G$. Let $L^{*}$ be the vertex set of the connected component of $G\left[K\cap\left(\bigcup_{H\in I_{\mathcal{H}}(L)}H\right)\right]$ containing $L$, and similarly, let $R^{*}$ be the vertex set of the connected component of $G\left[K\cap\left(\bigcup_{H\in I_{\mathcal{H}}(R)}H\right)\right]$ containing $R$. Clearly $I_{\mathcal{H}}(L^{*})$ and $I_{\mathcal{H}}(R^{*})$ both have weak diameter at most 5 in $\textbf{RIG}(G,\mathcal{H})$. See Figure˜3 for an illustration (in a case where $L^{*}$ and $R^{*}$ do not intersect). We split into two cases depending on whether or not $L^{*}$ and $R^{*}$ intersect.
Figure 3: An illustration of the constructed connected sets $L^{*}$ and RjRj (coloured purple) in a case where they do not intersect. Both $B_{1}$ and $B_{2}$ are coloured blue. The connected sets in $\mathcal{H}$ are coloured red or orange, with the orange ones being $H_{\ell},H_{\ell}^{\prime},H_{r},H_{r}^{\prime}$. The two orange paths indicate $L$ and $R$.

### v2

@@
String graphs are quasi-isometric to planar graphs
James Davies The author was supported by the Alexander von Humboldt Foundation in the framework of the Alexander von Humboldt Professorship of Daniel Kráľ endowed by the Federal Ministry of Education and Research.
Аннотация
@@
We further extend our techniques to show that every complete Riemannian surfaces $\Sigma$ of bounded Euler genus has a triangulation $G\subset\Sigma$ such that $G^{(1)}\hookrightarrow\Sigma$ is a quasi-isometry, where $G^{(1)}$ is the simplicial 1-skeleton of $G$.
@@ 1 Introduction
Every planar graph is a string graph, although the structure of string graphs is much more complicated. For instance, while planar graphs can be recognized in linear time [55], recognition of string graphs is NP-hard [62, 88]. There are at most $c^{n}n!$ labelled planar graphs on $n$ vertices [28, 79], while the number of labelled string graphs on $n$ vertices is $2^{(\frac{3}{4}+o(1)){n\choose 2}}$ [83]. Every planar graph is 4-colourable [6, 86], however there are triangle-free string graphs with arbitrarily large chromatic number [84].
@@
In other words, string graphs are quasi-isometric (or even bi-Lipschitz equivalent) to planar graphs. For finite string graphs the proof of Theorem˜1.1 yields a polynomial time (with respect to the size of a given representation of $S$ in the input) algorithm for generating such a planar graph $G$. Unfortunately, there are string graphs that require exponentially sized representations [61]. However, natural subclasses such as segment intersection graphs or more generally intersection graphs of convex compact path-connected subsets of $\mathbb{R}^{2}$, do have polynomially sized representations. We remark that if the string graph $S$ is locally finite, then the resulting planar graph $G$ is also clearly locally finite.
Theorem˜1.1 has numerous applications. We discuss some direct corollaries for string graphs. Afterwards, we shall discuss further quasi-isometry results on metric planar graphs and complete Riemannian surfaces before giving a couple more applications.
@@
One of the most natural classes of graphs to consider is of course planar graphs. Ostrovskii and Rosenthal [81] proved that planar graphs have bounded asymptotic dimension. Fujiwara and Papasoglu [42] improved this bound to 3 and Jørgensen and Lang [58] proved that planar graphs have asymptotic dimension and furthermore Assouad–Nagata dimension at most 2, which is tight due to the infinite planar grid. Asymptotic dimension (and Assouad–Nagata dimension in the case of graphs) are persevered by quasi-isometries. Thus, by Theorem˜1.1, we obtain the following.
@@
Bonamy, Bousquet, Esperet, Groenland, Liu, Pirot, and Scott [12] gave a proof that planar graphs have Assouad–Nagata dimension at most 2, which yields a polynomial time algorithm for computing the corresponding covers. Thus, the corresponding covers in Theorem˜1.2 can also be computed in polynomial time (in terms of the size of the provided representation of the string graph given in the input). For some algorithmic applications of such covers, see [8, 13, 14, 35, 37, 57, 63, 65].
@@
We now turn our attention to some further quasi-isometry results on metric planar graphs and complete Riemannian surfaces.
A metric graph $H$ is a graph along with an assignment of non-negative edge lengths (in this paper, the length of edges in (non-metric) graphs is otherwise always uniformly equal to 1). We can view metric graphs as a length space in which the edges are replaces by intervals of the corresponding lengths. When the edge lengths are within $D\subseteq(0,\infty)$, we say that $H$ is an $D$-metric graph. The coarse geometry of metric graphs tends to be harder to work with than that of non-metric (simplicial) graphs [64].
@@
Beyond the general interest in metric planar graphs and also in groups acting on planar surfaces [50], one other motivation of Georgakopoulos and Papasoglu [47] was their coarse Kuratowski conjecture that graphs forbidding $K_{5}$ and $K_{3,3}$ as $K$-fat minors are quasi-isometric to planar graphs. Although Theorem˜1.5 and Theorem˜1.1 provide substantial progress towards this conjecture, we remark that the closely related graph minor analogue that graphs forbidding $H$ as a $K$-fat minor are quasi-isometric to $H$-minor-free graphs is false [2, 26], (although it is known to be true for some small graphs $H$ [3, 5, 22, 43, 47, 69, 75]). Such $K$-fat $H$-minor-free graphs need not even be quasi-isometric to $H^{\prime}$-minor-free graphs for any graph $H^{\prime}$ [2].
It is well known (see [12, 18, 23, 48, 80, 87]) that complete Riemannian surfaces are quasi-isometric to locally finite [^2] A $(0,1]$-metric graph is locally finite if every bounded radius ball contains at most a finite number of vertices. $(0,1]$-metric graphs embeddable in the surface. This was used for instance to show that complete Riemannian surfaces of bounded Euler genus (such as complete Riemannian planes) have Assouad–Nagata dimension at most 2 [12] and in Maillot’s [68] aforementioned classification of finitely generated virtually free groups as being those quasi-isometric to certain planar Riemannian surfaces. Ntalampekos and Romney [80] also used a strong version for length surfaces as a step in their proof that any length surface is the Gromov–Hausdorff limit of polyhedral surfaces with controlled geometry. They gave further applications of this, including a new proof of the Bonk-Kleiner theorem [15] characterizing Ahlfors 2-regular quasispheres.
By Theorem˜1.5, we obtain the following refinement, which significantly generalizes the previously discussed fact that finitely generated groups quasi-isometric to complete Riemannian planes are quasi-isometric to planar graphs.
@@
In fact, we shall further extend our techniques to refine Theorem˜1.6 in three ways. Firstly, we obtain a quasi-isometry which is also an embedding of the graph into the Riemannian plane (as in [12, 18, 23, 48, 80, 87]). Secondly, we also extend this result to Riemannian surfaces of bounded Euler genus. Lastly, we show that such a graph and embedding can be chosen to be a triangulation. A key ingredient in this last refinement is a recent result of Georgakopoulos and Vigolo [48] that complete Riemannian surfaces admit triangulations by metric graphs that are quasi-isometric to the surface. For an embedded graph $G\subset\Sigma$, we denote by $G^{(1)}$ its simplicial 1-skeleton (in which all edges have length 1).
Theorem 1.7.
Let $\Sigma$ be a complete Riemannian surface (without boundary) of bounded Euler genus. Then $\Sigma$ has a triangulation $G\subset\Sigma$ such that $G^{(1)}\hookrightarrow\Sigma$ is a quasi-isometry.
The quasi-isometry in Theorem˜1.7 can be taken to be a $(10^{6},10^{6}+10^{4}g)$-quasi-isometry where $g$ is the Euler genus of the Riemannian surface. The dependence on the Euler genus is necessary since for each fixed $M\geqslant 1$, there are complete Riemannian surfaces of genus at most $g$ that admit no such $(M,\Omega(\sqrt{\log g}))$-quasi-isometric simplicial triangulation [27]. In particular, Theorem˜1.7 does not hold for complete Riemannian surfaces of unbounded genus [27]. Georgakopoulos and Vigolo [48] recently showed that complete Riemannian surfaces with uniform nets also admit quasi-isometric simplicial triangulations as in Theorem˜1.7, and Bowditch [17] showed that complete Riemannian manifolds of bounded geometry also admit such simplicial triangulations.
We remark that while we prove Theorem˜1.7 for Riemannian surfaces without boundary, it is possible to extend Theorem˜1.7 to Riemannian surfaces with boundary with more technical arguments and with the same resulting $(10^{6},10^{6}+10^{4}g)$-quasi-isometry bounds.
We expect our quasi-isometry results to have many further applications. We quickly give two more concerning $\ell_{1}$-embeddings and the coarse Menger conjecture.
Theorem 1.8.
For a more extensive survey and discussion of the applications of such embeddings, see [52].
Albrechtsen, Huynh, Jacobs, Knappe and Wollan [4], and independently Georgakopoulos and Papasoglu [47] conjectured a coarse analogue of Menger’s [72] theorem that for a graph $G$ containing two vertex sets $S,T$, either there are $k+1$ paths between $S$ and $T$ pairwise at distance at least $c$, or there is a set of at most $k$ balls in $G$ of radius at most $f(c,k)$ such that every path between $S$ and $T$ intersects one of these balls. This turns out to be false in general [74], and even in the weak version where up to $f(c,k)$ balls are allowed [76]. However, it is true for certain classes such as those of bounded path-width [77]. A particularly interesting open case is that of planar graphs. Nguyen, Scott, and Seymour [78] proved a version for paths between vertices on the outerface of a planar graph. By Theorem˜1.5, we obtain the following extension of their result to metric planar graphs [^3] We remark that there is a technicality of preserving vertices on the outerface, however, the proof of Theorem 1.5 gives this. (they proved the same theorem for (simplicial) graphs, just with a much better constant than $10^{10}$).
Theorem 1.9.
Let $c,k\geqslant 0$, and let $G$ be a metric planar graph, with vertex sets $S,T$ on the outerface. Then either:
  • there are $k+1$ paths between $S$ and $T$, pairwise at distance least $c$; or
  • there is a set of at most $k$ balls in $G$ of radius at most $10^{10}k^{3}c$ of G, such that every path between $S$ and $T$ intersects one of these balls.
One can also obtain versions of Theorem˜1.9 for string graphs and complete Riemannian surfaces homeomorphic to a disk.
@@
In Section˜2 we introduce basic definitions, deduce Theorem˜1.5 from Theorem˜1.1, and also reduce proving Theorem˜1.1 to its finite case (Theorem˜2.5). We also introduce the necessary framework for our approach to proving Theorem˜1.1, namely through viewing string graphs as region intersection graphs, and the notion of ‘‘impressions’’ which provides a convenient way of finding quasi-isometries from region intersection graphs. The rest of the proof in Sections˜3, 4 and 5 is essentially a delicate technical induction. In Section˜3 we prove a weakening of Theorem˜1.1 for the subclass outerstring graphs (Theorem˜3.1). This is used in Section˜4 to find a technical structure tailored for use in the inductive argument that we call an ‘‘encasing’’ (see Lemma˜4.4). Finally, in Section˜5, we prove Theorem˜1.1 using Lemma˜4.4. Lastly, in Section˜6, we build on one version of our main technical result (Theorem˜5.2) to prove Theorem˜1.7.
Remark 1.10.
More recently, Chang, Conroy, Tan, and Zheng [21] have independently obtained another proof of a result very similar to Theorem˜1.1. While our proof is self contained and builds the quasi-isometry from first principles, theirs adapts a shortcut partition construction using gridtrees developed in [20]. We remark that their motivation and applications differs significantly from the present paper and we encourage the interested reader to also see [21] for further applications. For instance, they construct compact (approximate) distance oracles for string graphs and also prove that for any $\epsilon>0$, $n$-vertex string graphs have $(1+\epsilon,O(\epsilon^{-4}\log^{18}n))$-quasi-isometric metric planar graphs.
In [24], we will build on the work of the present paper and [21] to prove versions of Theorem˜1.5 and Theorem˜1.1 for minor-free graphs.
@@ 2 Preliminaries
For $M,A\geqslant 0$ with $M\geqslant 1$, an $(M,A)$-quasi-isometry from a metric space $G$ to another metric space $H$ is a map $f:G\to H$ such that
  1. 1.
    $M^{-1}d_{G}(u,v)-A\leqslant d_{H}(f(u),f(v))\leqslant Md_{G}(u,v)+A$ for every $u,v\in G$, and
  2. 2.
    for every $h\in H$, there exists some $v\in G$ with $d_{H}(h,f(v))\leqslant A$.
We say that $G$ is $(M,A)$-quasi-isometric to $H$ if there exists a $(M,A)$-quasi-isometry from $G$ to $H$. Two metric spaces $G$ and $H$ are quasi-isometric if they are $(M,A)$-quasi-isometry for some $M,A\geqslant 0$. Quasi-isometries for graphs are of course defined in the same expected way.
@@
Furthermore, $f$ can be chosen to be any function $f:\mathcal{H}\to\mathcal{M}$ such that $f(H)\in I_{\mathcal{M}}(H)$ for all $H\in\mathcal{H}$.
Доказательство.
Let $f:\mathcal{H}\to\mathcal{M}$ be chosen so that $f(H)\in I_{\mathcal{M}}(H)$ for all $H\in\mathcal{H}$. Clearly such a function exists since $I_{\mathcal{M}}(H)$ is always non-empty.
@@
Next we show that Theorem˜1.1 implies Theorem˜1.5. This is immediate from the following lemma.
@@
Доказательство.
Consider $H$ embedded in the plane. For each $u\in V(H)$, let $S_{u}$ be the subset of the plane consisting of all points of $H$ at distance at most 1 in $H$ from $u$. Let $S$ be the string graph with vertex set $V(H)$ in which each $u\in V(S)=V(H)$ is represented by $S_{u}$. Consider $u,v\in V(H)$ and let $P$ be a geodesic path from $u$ to $v$ in $H$. Choose vertices $x_{0},\ldots,x_{t}$ of $H$ along $P$ so that $x_{0}=u$, $x_{t}=v$ and for each $0\leqslant i<t$, the distance along $P$ between $x_{i}$ and $x_{i+1}$ is in $[1,2]$ (or if $d_{H}(u,v)<1$, we just take $x_{0}=u$, $x_{1}=v$). So, $\lceil\frac{1}{2}d_{H}(u,v)\rceil\leqslant t\leqslant\lfloor d_{H}(u,v)\rfloor$. The lemma now follows since $x_{i}$ is adjacent to $x_{i+1}$ in $S$ for each $0\leqslant i<t$. ∎
@@ 3 Outerstring graphs
Outerstring graphs are equivalent to region intersection graphs $\textbf{RIG}(G,\mathcal{H})$ for which $G$ is planar and every $H\in\mathcal{H}$ contains a vertex of the outer face of $G$. They can be thought of as an ‘‘outerplanar’’ analogue of string graphs, and are thus much simpler. The main aim of this section is to prove the following theorem on finding certain restricted impressions of such region intersection graphs. This will be crucial to our arguments on string graphs (or strictly speaking, Lemma˜3.4 and Lemma˜3.2 which easilly imply the following theorem will be).
@@
Рис. 1: The graph $F_{6}$.
By Proposition˜2.1, Theorem˜3.1 implies that outerstring graphs are quasi-isometric to outerplanar graphs. This result is already easily implied by a theorem that graphs forbidding $K_{2,3}$ as a fat minor are quasi-isometric to cacti [22, 43], where a cacti is a planar graph in which every edge is on the outer face, or equivalently, a graph in which each maximal 2-connected subgraph is either an edge of a cycle. One might hope that Theorem˜3.1 could be improved so that $\textbf{IM}(G,\mathcal{M})$ is also a cacti. However this is not possible. For every $\ell\geqslant 3$, let $F_{\ell}$ be the outerplanar graph consisting of a cycle of length $\ell$ where for every edge of this cycle, we identify the edge with an edge of another cycle of length $\ell$ (see Fig.˜1 for an illustration of the graph $F_{6}$). Then it can easily be seen that there is no $x,y$ such that every $(F_{\ell},\{\{u,v\}:uv\in E(F_{\ell})\})$ has a $(x,y)$-impression $(F_{\ell},\mathcal{M}_{\ell})$ in which $\textbf{IM}(F_{\ell},\mathcal{M}_{\ell})$ is a cacti. It is however a nice exercise to show that outerplanar graphs are quasi-isometric to cacti. With this, Theorem˜3.1 and Proposition˜2.1 then yeild another proof that outerstring graphs are quasi-isometric to cacti.
@@
Доказательство.
Clearly it is enough to handle the case that $G$ is connected. Choose some $H_{0}\in\mathcal{H}$ and for every non-negative integer $i$, let $L_{i}$ be the vertices of $\textbf{RIG}(G,\mathcal{H})$ at distance $i$ from $H_{0}$ in $\textbf{RIG}(G,\mathcal{H})$. For each $i$ we define the following two vertex sets in $G$. Let $D_{i}=\bigcup_{H\in L_{i}}H$, and let $A_{i}=D_{i}\backslash D_{i-1}$ (where $D_{-1}=\emptyset$). For each $A_{i}$, let $\mathcal{M}_{i}$ be the maximal connected vertex sets of $G[A_{i}]$. Let $\mathcal{M}=\bigcup_{i=0}^{\infty}\mathcal{M}_{i}$. See Fig.˜2 for an illustration. We must show that $(G,\mathcal{M})$ is a $(11,9)$-impression of $(G,\mathcal{H})$.
Рис. 2: An illustration of the resulting collection of disjoint connected sets $\mathcal{M}$. The connected sets in $\mathcal{H}$ are the red sets and the sets in $\mathcal{M}$ are the blue ones. The dashed purple lines indicate the partition of $\mathcal{M}$ into $\bigcup_{i=0}^{\infty}\mathcal{M}_{i}$.
@@
Lastly, we can now assume that $K$ satisfies Item˜4. Let $B_{1},B_{2},F$ be as in Item˜4. Note that the intersection of $K$ with the vertices of the outer face of $G$ induces a connected graph since $B_{1}$ and $B_{2}$ both contain a vertex on the outer face of $G$ and the face $F$ (which is distinct from the outer face). Let $Q$ be the vertex set of this connected induced subgraph. Let $p_{1}p_{2}\ldots p_{t}$ be the walk in $G[Q]$ along the outer face of $G$ (and containing all of) $Q$, where $p_{1}$ touches $B_{1}$ and $p_{t}$ touches $B_{2}$. Choose $1\leqslant\ell\leqslant t$ maximum such that that there exists a path with vertex set $L$ between $p_{\ell}$ and $N_{G}(B_{1})$ such that $L$ is contained in $K\cap(H_{\ell}\cup H^{\prime}_{\ell})$ for some intersecting (but not necessarily distinct) $H_{\ell},H^{\prime}_{\ell}\in\mathcal{H}$. Similarly, choose $1\leqslant r\leqslant t$ minimum such that that there exists a path with vertex set $R$ between $p_{r}$ and $N_{G}(B_{2})$ such that $R$ is contained in $K\cap(H_{r}\cup H^{\prime}_{r})$ for some intersecting (but not necessarily distinct) $H_{r},H^{\prime}_{r}\in\mathcal{H}$. Note that such $\ell,r$ exist since $K$ touches exactly $B_{1},B_{2}$ from $\mathcal{B}$, and $K$ (and every $H\in\mathcal{H}$) contains a vertex incident to the outer face of $G$. Let $L^{*}$ be the vertex set of the connected component of $G\left[K\cap\left(\bigcup_{H\in I_{\mathcal{H}}(L)}H\right)\right]$ containing $L$, and similarly, let $R^{*}$ be the vertex set of the connected component of $G\left[K\cap\left(\bigcup_{H\in I_{\mathcal{H}}(R)}H\right)\right]$ containing $R$. Clearly $I_{\mathcal{H}}(L^{*})$ and $I_{\mathcal{H}}(R^{*})$ both have weak diameter at most 5 in $\textbf{RIG}(G,\mathcal{H})$. See Fig.˜3 for an illustration (in a case where $L^{*}$ and $R^{*}$ do not intersect). We split into two cases depending on whether or not $L^{*}$ and $R^{*}$ intersect.
Рис. 3: An illustration of the constructed connected sets $L^{*}$ and $R^{*}$ (coloured purple) in a case where they do not intersect. Both $B_{1}$ and $B_{2}$ are coloured blue. The connected sets in $\mathcal{H}$ are coloured red or orange, with the orange ones being $H_{\ell},H_{\ell}^{\prime},H_{r},H_{r}^{\prime}$. The two orange paths indicate $L$ and $R$.